Hydrogen (a moles)  and iodine (b moles) react to give 2x moles of the HI at equilibrium.  The total number of moles at equilibrium is 

Grade:11

1 Answers

Khimraj
5 years ago
H2    +   I2    = 2HI
a          b          0
a-x       b-x       2x
total moles at equilibrium = a-x + b-x + 2x = a+b
Hope it clears................................